# A LLM-Based Ranking Method for the Evaluation of Automatic Counter-Narrative Generation

### v1

@@ Abstract
The proliferation of misinformation and harmful narratives in online discourse has underscored the critical need for effective Counter Narrative (CN) generation techniques. However, existing automatic evaluation methods often lack interpretability and fail to capture the nuanced relationship between generated CNs and human perception. Aiming to achieve a higher correlation with human judgments, this paper proposes a novel approach to asses generated CNs that consists on the use of a Large Language Model (LLM) as a evaluator. By comparing generated CNs pairwise in a tournament-style format, we establish a model ranking pipeline that achieves a correlation of $0.88$ with human preference. As an additional contribution, we leverage LLMs as zero-shot (ZS) CN generators and conduct a comparative analysis of chat, instruct, and base models, exploring their respective strengths and limitations. Through meticulous evaluation, including fine-tuning experiments, we elucidate the differences in performance and responsiveness to domain-specific data. We conclude that chat-aligned models in ZS are the best option for carrying out the task, provided they do not refuse to generate an answer due to security concerns.
@@ 1 Introduction
The proliferation of misinformation and the dissemination of harmful narratives has underscored the urgent need for effective strategies to combat Hate Speech (HS). This necessity has drawn significant attention to the field of automatic CN generation, where considerable research has been conducted on the use of LLMs to fulfill this task with impressive results Chung et al. (2021); Tekiroğlu et al. (2022). However, difficulties in automatically assessing the quality of the generated CNs remain. As is common in text generation tasks, while manual evaluation is expensive, time-consuming, and subjective, existing automatic methods often fail to provide comprehensive insights or capture the nuanced relationship between generated text and human perception, overlooking crucial aspects of effectiveness and relevance (Ni’mah et al., 2023). Finally, the evaluation problem is exacerbated by the lack of a ’universal truth’ and the significant variations among possible references (see example in Table 1).
@@
Table 1: Sample set of proposed CNs for a single HS instance from the CONAN corpus. Although only two references are shown in the table, the corpus includes a total of 36 candidate CNs as the gold standard for the presented instance of HS.
In this paper we address these shortcomings by presenting a novel approach to automatically evaluate CNs, aiming for a higher correlation to human preference than the traditional metrics commonly used to evaluate natural language generation (NLG) tasks. We propose to evaluate generated CNs pairwise in a tournament-style format and determine the outcome of said tournaments without the need for human intervention through the use of a Judge Model explicitly trained to assess the quality of generated text. The outcomes of the tournaments will be used to rank models according to their performance. This approach allows us to decompose a subjective task like CN evaluation into a series of simpler binary classification problems.
As an additional contribution, we focus on evaluating the inherent ability of LLMs as ZS CN generators. Leveraging state-of-the-art open-source LLMs, we seek to explore their potential in generating CNs that effectively challenge and mitigate the influence of misinformation and harmful narratives. We examine three variants within the same model family: base, instruction-tuned, and chat-aligned. This enables us to inspect their unique strengths and limitations to determine the optimal choice for the task. Finally, we fine-tune the models on HS-CN pair data to compare their performance against ZS performance, assessing whether fine-tuning offers any significant improvement in our scenario. We conclude that chat-aligned models in a ZS setting are the best option for carrying out the task, provided they do not refuse to generate an answer due to security concerns [^1] Chat-aligned models, designed to adhere to safety and ethical guidelines, may sometimes decline to respond to certain prompts. This refusal is typically in place to prevent the generation of harmful, inappropriate, or sensitive content.. Code will be made publicly available upon publication.
@@ 2 Related Work
Several datasets have been introduced to aid in the advancement of CN generation. The first large-scale, multilingual, expert-based dataset, Counter Narratives through Nichesourcing (CONAN) (Chung et al., 2019), consists of HS-CN pairs in English, French, and Italian, focusing exclusively on Islamophobia. This corpus served as the foundation for the development of MultiTarget CONAN (MT-CONAN) (Fanton et al., 2021a), which includes 8 hate-speech targets such as women and individuals with disabilities. Additionally, the DIALOCONAN dataset (Bonaldi et al., 2022), which contains fictitious dialogues between a hater and a Non-Governmental Organization (NGO) operator, and the Knowledge-grounded Hate Countering dataset (Chung et al., 2021), featuring HS-CN pairs with the background knowledge used for constructing the CNs have been introduced. Some work in adapting these corpora to other languages has also been done, such as CONAN-EUS Bengoetxea et al. (2024), a Basque and Spanish translation of the original CONAN dataset, and CONAN-MT-SP (Vallecillo Rodríguez et al., 2024), a Spanish version of MT-CONAN.
Assessing the impact and effectiveness of the generated CNs remains a crucial aspect of this research domain. Evaluating CN is particularly challenging because there are many acceptable answers to a given HS, and is often very difficult to assess what constitutes a good answer. Evaluation is usually done either through automatic or manual methods. Automatic methods involve the use of metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and BERTScore Zhang et al. (2020), which are standard evaluation metrics in tasks such as Machine Translation or Text Summarization. However, these metrics are known to be often weakly correlated with human judgment Sai et al. (2022), particularly on tasks that require creativity Ni’mah et al. (2023). Other automatic metrics that have been used for NLG evaluation including Repetition Rate (RR) (Bertoldi et al., 2013), which measures diversity in the generated answers, or Novelty Wang and Wan (2018), which encourages the model to generate answers that differ from the text in the training data. We must note that none of the aforementioned metrics take the input HS into account, missing a crucial aspect in the evaluation of CNs. As far as we know, no previous work has analyzed the correlation of these metrics with human judgment for CN evaluation.
Due to the limitations of automatic metrics, final assessments frequently rely on manual evaluations. Nonetheless, manual evaluation is often a costly process, and finding evaluators with adequate task comprehension can be challenging. Moreover, the subjective nature of the task adds another layer of complexity to the evaluation process. To mitigate this subjectivity, various efforts have been made to identify key aspects that assess the quality of CNs. Unfortunately, consensus on these key aspects is still lacking, as different authors consider factors such as relatedness, specificity, richness, coherence, grammaticality, suitableness, informativeness, diversity, relevance, language quality, offensiveness or stance Chung et al. (2021); Ashida and Komachi (2022); Bengoetxea et al. (2024).
Recently, to address the limitations of traditional metrics and manual evaluation, LLMs are being employed to directly assess the quality of generated text. Leveraging LLMs to measure NLG has been shown to exhibit stronger correlation with human assessment compared to conventional reference-based evaluation Nimah et al. (2023); Chiang and Lee (2023); Wang et al. (2023); Liu et al. (2023), and it remains an efficient and automated approach. Some works include using commercial models such as GPT-4 Wang et al. (2023); Liu et al. (2023), while others focus on training specialized models for evaluation tasks, resulting in tools like PandaLM (Wang et al., 2024), JudgeLM Zhu et al. (2023), and UniEval (Zhong et al., 2022). These LLMs can be either used to ascertain quantitative aspects to measure the quality of the generated text (Zhong et al., 2022; Ke et al., 2022), or to show preferences between two texts generated by different systems.
There are few works on automatic CN evaluation. Contemporary to our work, Jones et al. (2024) use LLMs to evaluate CNs based on five different aspects considered relevant for their effectiveness in combating HS. In a strategic shift, we propose to use LLMs to compare outputs from different models between them and ultimately obtain a ranking of the best models for CN generation that correlates with human judgments.
3 Methodology
3.1 Models
In this study, we used auto-regressive models for CN generation. Specifically, we work with three variants of the Mistral model family (Jiang et al., 2023) as well as the Llama 2 Chat model (Touvron et al., 2023). The Mistral variants include the Mistral base model, the Mistral-Instruct model, and Zephyr (Tunstall et al., 2023), which is a chat-aligned model based on Mistral. The selection of these three variants enables the comparison of chat-aligned, instruction-tuned, and base model performance and behavior. Llama 2 was selected as to compare the results with Mistral models due to its relevance in the field and potential for providing complementary insights into CN generation. All models are 7B parameter models, consistent with the available Mistral model size, to ensure comparability of results. Additionally, the weights of all the models used in this study are publicly available. The specific versions of the employed models are listed in Table 3.
3.2 Corpus
@@
In order to test the generalizability of our method, we conduct the analysis on two distinct datasets: Counter Narratives through Nichesourcing (CONAN) (Chung et al., 2019) and Multi-Target CONAN (MT-CONAN) (Fanton et al., 2021b). Corpus statistics are presented in Table 2.
@@ 3.3 Evaluation Metrics
Based on previous work on CN generation, we opted to use BLEU, ROUGE-L and BERTScore. BLEU is a precision-based metric that measures the similarity between a candidate text and one or more reference texts. It computes the geometric mean of modified n-gram precision and applies a brevity penalty to discourage short translations. BLEU is widely used in machine translation tasks. ROUGE-L, on the other hand, focuses on the recall of content units. It calculates the longest common subsequence between the candidate sequence and the reference sequence, normalizing by the length of the reference sequence. ROUGE-L is commonly employed in text summarization tasks. BERTScore leverages contextual embeddings from pre-trained BERT models to compute the similarity between candidate and reference sentences. It computes the score based on the cosine similarity between BERT embeddings, providing a measure of semantic similarity. BERTScore has demonstrated effectiveness across various NLG tasks, including machine translation and text summarization.
@@
For reference-free metrics, we opted to use Repetition Rate (RR) (Bertoldi et al., 2013), which is computed by calculating the non-singleton n-grams that are repeated in the generated text (Bertoldi et al., 2013) and Novelty Wang and Wan (2018) that is computed by calculating the non-singleton n-grams from the generated text that appear in the train data. While RR aims to capture the diversity in the generated text, Novelty measures how different the generated text is from the training data. It should be noted that Novelty is less valuable when evaluating models that were used in a ZS setting, as there is no training involved.
@@
Finally, we consider the use of JudgeLM as an evaluator. JudgeLM is a scalable Judge Model based on Vicuna that was designed to evaluate LLMs in open-ended scenarios. It was trained using a large-scale dataset consisting on LLM-generated answers for diverse NLG tasks and detailed judgments from GPT-4. Remarkably, it achieves an agreement rate exceeding 90% in some tasks, surpassing even human-to-human agreement levels Zhu et al. (2023). While JudgeLM supports different evaluation methods, such as comparing single answers against a reference or comparing multiple answers simultaneously, we decided to use it to compare generated CNs pairwise, as described in Section 4.1. This eliminates the problem of needing a reliable reference and instead focuses on determining which of the available options is the best, simplifying the task. Comparing the CNs against each other also avoids the ambiguity of the open-ended scenario we would face if we decided to evaluate them individually. We must note that, as seen in Appendix C.5, where the prompting format of the model is shown, JudgeLM considers the instance of HS when comparing the effectiveness of CNs, evaluating their effectiveness in relation to a specific instance of HS, not just as standalone sentences.
JudgeLM operates in two modes: fast evaluation activated or deactivated. In fast evaluation mode, the model outputs two scores, one for each CN, providing an overall assessment of their value. When deactivated, the model supplements these scores with arguments explaining the rationale behind them. Both during the development stage and while conducting the result analysis, we deactivated fast mode to ensure a comprehensive evaluation. However, for creating the ranks, we opted to deactivate the argumentation feature and solely output the scores. This decision was made because generating arguments significantly increases inference time and argumentation was deemed unnecessary for our specific task.
@@ 4 Evaluation Framework
In this section we present a cost-effective pairwise rank-based evaluation paradigm designed to assess the performance of CN generation systems in alignment with human preference. The method is detailed in Section 4.1. In addition, in Section 4.2, we describe the additional manual evaluation conducted to provide a detailed assessment according to various relevant aspects contributing to the effectiveness of a CN, as presented in Bengoetxea et al. (2024).
@@ 4.1 Pairwise Rank-Based Evaluation
We propose an "A vs B" comparative setup to rank models with respect to their CN generation skills. Suppose we have $n$ models to rank and we want to evaluate their performance in a test set consisting of $h$ HS instances. First, each one of the $n$ models will generate a CN for each instance of HS in the test set. After that, the generated CNs will be pitted against each other in "A vs B" tournaments. In the end, we are left with (${n\choose 2}\cdot h=\frac{n!}{2!(n-2)!}\cdot h$ tournaments. We will use either automatic or manual evaluation to decide the outcome of said tournaments. Based on the results, each of the $n$ models will receive a performance score using the following point scheme: the winning model receives $1$ point, the losing model receives 00 points, and in the case of a tie, both models receive $0.5$ points. Finally, the models will be ranked based on the obtained score.
@@
For automatic evaluation we will prompt JudgeLM to output 2 scores: one for each proposed CN. The winner will be determined by comparing these scores: the CN with the highest score is the winner. If both scores are the same, it results in a tie.

### v3

@@ Abstract
This paper proposes a novel approach to evaluate Counter Narrative (CN) generation using a Large Language Model (LLM) as an evaluator. We show that traditional automatic metrics correlate poorly with human judgements and fail to capture the nuanced relationship between generated CNs and human perception. To alleviate this, we introduce a model ranking pipeline based on pairwise comparisons of generated CNs from different models, organized in a tournament-style format. The proposed evaluation method achieves a high correlation with human preference, with a $\rho$ score of $0.88$. As an additional contribution, we leverage LLMs as zero-shot CN generators and provide a comparative analysis of chat, instruct, and base models, exploring their respective strengths and limitations. Through meticulous evaluation, including fine-tuning experiments, we elucidate the differences in performance and responsiveness to domain-specific data. We conclude that chat-aligned models in zero-shot are the best option for carrying out the task, provided they do not refuse to generate an answer due to security concerns.
@@ 1 Introduction
The proliferation of misinformation and the dissemination of harmful narratives has stressed the urgent need for effective strategies to combat Hate Speech (HS). This necessity has drawn significant attention to the field of automatic CN generation, where considerable research has focused on the use of LLMs to fulfill this task (Chung et al., 2021; Tekiroğlu et al., 2022). However, difficulties in automatically assessing the quality of the generated CNs remain. As is common in text generation tasks, while manual evaluation is expensive, time-consuming, and subjective, existing automatic methods often fail to provide comprehensive insights or capture the nuanced relationship between generated text and human perception, overlooking crucial aspects of effectiveness and relevance (Ni’mah et al., 2023). Finally, the problem of CN evaluation is exacerbated by the lack of a ’universal truth’ and the significant variations among possible references, as shown in Table 1.
In this paper we address the limitations of traditional evaluation metrics for CN generation by proposing a novel automatic evaluation approach. This method is motivated by the need to improve upon existing metrics like BLEU, ROUGE and BERTScore (Papineni et al., 2002; Lin, 2004; Zhang et al., 2020) which do not consider the specific HS to which the CN is responding to, an essential aspect for accurately assessing the quality of CNs. Our approach evaluates generated CNs pairwise in a tournament-style format, with outcomes determined without human intervention using JudgeLM (Zhu et al., 2023), a model explicitly trained to assess the quality of text. Thus, JudgeLM enables pairwise comparisons of CNs, addressing the subjectivity of the task by breaking it down into simpler binary classification problems. To ensure the generalizability of our approach for the task, we test it on two distinct corpora: CONAN (Chung et al., 2019) and CONAN-MT (Fanton et al., 2021a). Using various models to generate texts of different quality means that we can evaluate model performance across a spectrum of CN quality, ensuring that even subtle distinctions between good texts can be captured. This approach ultimately aims for a higher correlation to human preference than traditional metrics and ranks models based on their ability to generate high-quality CNs.
As an additional contribution, we evaluate the inherent ability of LLMs as zero shot (ZS) CN generators. By leveraging state-of-the-art open-source LLMs we seek to explore their potential in generating CNs that effectively challenge and mitigate the influence of misinformation and harmful narratives. We examine three variants within the same model family: base, instruction-tuned, and chat-aligned. This enables us to inspect their unique strengths and limitations to determine the optimal choice for the task. Finally, we fine-tune the models on HS-CN pair data to compare their performance against ZS performance, assessing whether fine-tuning offers any significant improvement in our scenario. We conclude that chat-aligned models in a ZS setting are the best option for carrying out the task, provided they do not refuse to generate an answer due to security concerns [^1] Chat-aligned models, designed to adhere to safety and ethical guidelines, may sometimes decline to respond to certain prompts. This refusal is typically in place to prevent the generation of harmful, inappropriate, or sensitive content.. Code is available at https://github.com/hitz-zentroa/cn-eval.
@@
Table 1: Example of a HS and two reference CNs, taken from the CONAN corpus. The CNs express diverse points, and show the variability and diversity of arguments used to combat HS.
@@ 2 Related Work
Several datasets have been introduced to aid in the advancement of CN generation. The first large-scale, multilingual, expert-based dataset, Counter Narratives through Nichesourcing (CONAN) (Chung et al., 2019), consists of HS-CN pairs in English, French, and Italian, focusing exclusively on Islamophobia. This corpus served as the foundation for the development of MultiTarget CONAN (MT-CONAN) (Fanton et al., 2021a), which includes 8 hate-speech targets such as women and individuals with disabilities. Additionally, the DIALOCONAN dataset (Bonaldi et al., 2022), which contains fictitious dialogues between a hater and a Non-Governmental Organization (NGO) operator, and the Knowledge-grounded Hate Countering dataset (Chung et al., 2021), featuring HS-CN pairs with the background knowledge used for constructing the CNs have been introduced. Some work in adapting these corpora to other languages has also been done, such as CONAN-EUS (Bengoetxea et al., 2024), a Basque and Spanish translation of the original CONAN dataset, and CONAN-MT-SP (Vallecillo Rodríguez et al., 2024), a Spanish version of MT-CONAN.
@@
Assessing the impact and effectiveness of the generated CNs remains a crucial aspect of this research domain. Evaluating CN is particularly challenging because there are many acceptable answers to a given HS, and is often very difficult to assess what constitutes a good answer. Evaluation is usually done either through automatic or manual methods. Automatic methods involve the use of metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang et al., 2020), which are standard evaluation metrics in tasks such as Machine Translation or Text Summarization. However, these metrics are known to be often weakly correlated with human judgment (Sai et al., 2022), particularly on tasks that require creativity (Ni’mah et al., 2023). Other automatic metrics that have been used for NLG evaluation including Repetition Rate (RR) (Bertoldi et al., 2013), which measures diversity in the generated answers, or Novelty (Wang and Wan, 2018), which encourages the model to generate answers that differ from the text in the training data. We must note that none of the aforementioned metrics take the input HS into account, missing a crucial aspect in the evaluation of CNs. As far as we know, no previous work has analyzed the correlation of these metrics with human judgment for CN evaluation.
Due to the limitations of automatic metrics, final assessments frequently rely on manual evaluations. Nonetheless, manual evaluation is often a costly process, and finding evaluators with adequate task comprehension can be challenging. Moreover, the subjective nature of the task adds another layer of complexity to the evaluation process. To mitigate this subjectivity, various efforts have been made to identify key aspects that assess the quality of CNs. Unfortunately, consensus on these key aspects is still lacking, as different authors consider factors such as relatedness, specificity, richness, coherence, grammaticality, suitableness, informativeness, diversity, relevance, language quality, offensiveness or stance (Chung et al., 2021; Ashida and Komachi, 2022; Bengoetxea et al., 2024).
Recently, to address the limitations of traditional metrics and manual evaluation, LLMs are being employed to directly assess the quality of generated text. Leveraging LLMs to measure NLG has been shown to exhibit stronger correlation with human assessment compared to conventional reference-based evaluation (Nimah et al., 2023; Chiang and Lee, 2023; Wang et al., 2023; Liu et al., 2023), and it remains an efficient and automated approach. Some works include using commercial models such as GPT-4 (Wang et al., 2023; Liu et al., 2023), while others focus on training specialized models for evaluation tasks, resulting in tools like PandaLM (Wang et al., 2024), JudgeLM (Zhu et al., 2023), and UniEval (Zhong et al., 2022). These LLMs can be either used to ascertain quantitative aspects to measure the quality of the generated text (Zhong et al., 2022; Ke et al., 2022), or to show preferences between two texts generated by different systems.
There are few works on automatic CN evaluation. Contemporary to our work, (Jones et al., 2024) use LLMs to evaluate CNs based on five different aspects considered relevant for their effectiveness in combating HS. In a strategic shift, we propose to use LLMs to compare outputs from different models between them and ultimately obtain a ranking of the best models for CN generation that correlates with human judgments.
3 Methodology
3.1 Models
We use publicly available auto-regressive models for CN generation. Specifically, we work with three variants of the Mistral model family (Jiang et al., 2023) as well as the Llama 2 Chat model (Touvron et al., 2023). The Mistral variants include the Mistral base model, the Mistral-Instruct model, and Zephyr (Tunstall et al., 2023), which is a chat-aligned model based on Mistral. The selection of these three variants enables the comparison of chat-aligned, instruction-tuned, and base model performance and behavior. Llama 2 was selected as to compare the results with Mistral models due to its relevance in the field and potential for providing complementary insights into CN generation. All models are 7B parameter models, consistent with the available Mistral model size, to ensure comparability of results. The specific versions of the employed models are listed in Table 3.
3.2 Corpus
Model Version Type lrCONAN lrMT-CONAN
mistral v0.1 Base 1e-5 1e-4
mistral-instruct v0.2 Instruct 6e-6 3e-5
zephyr Beta Chat 6e-6 1e-4
llama-chat llama 2 Chat 2e-5 1e-3
In order to test the generalizability of our method, we experiment on two distinct datasets: Counter Narratives through Nichesourcing (CONAN) (Chung et al., 2019) and Multi-Target CONAN (MT-CONAN) (Fanton et al., 2021b). Corpus statistics are presented in Table 2.
@@ 3.3 Evaluation Metrics
Based on previous work on CN generation, we opted to use BLEU, ROUGE-L and BERTScore. BLEU is a precision-based metric that measures the similarity between a candidate text and one or more reference texts. It computes the geometric mean of modified n-gram precision and applies a brevity penalty to discourage short translations. BLEU is widely used in machine translation tasks. ROUGE-L, on the other hand, focuses on the recall of content units. It calculates the longest common subsequence between the candidate sequence and the reference sequence, normalizing by the length of the reference sequence. ROUGE-L is commonly employed in text summarization tasks. BERTScore leverages contextual embeddings from pre-trained BERT models to compute the similarity between candidate and reference sentences. It computes the score based on the cosine similarity between BERT embeddings, providing a measure of semantic similarity. BERTScore has demonstrated effectiveness across various NLG tasks, including machine translation and text summarization. We report BERTScore F1 in our experiments.
@@
For reference-free metrics, we opted to use Repetition Rate (RR) (Bertoldi et al., 2013), which is computed by calculating the non-singleton n-grams that are repeated in the generated text (Bertoldi et al., 2013) and Novelty (Wang and Wan, 2018) that is computed by calculating the non-singleton n-grams from the generated text that appear in the train data. While RR aims to capture the diversity in the generated text, Novelty measures how different the generated text is from the training data. It should be noted that Novelty is less valuable when evaluating models that were used in a ZS setting, as there is no training involved.
@@
Finally, we consider the use of JudgeLM as an evaluator. JudgeLM is a scalable Judge Model based on Vicuna that was designed to evaluate LLMs in open-ended scenarios. It was trained using a large-scale dataset consisting on LLM-generated answers for diverse NLG tasks and detailed judgments from GPT-4. Remarkably, it achieves an agreement rate exceeding 90% in some tasks, surpassing even human-to-human agreement levels (Zhu et al., 2023). While JudgeLM supports different evaluation methods, such as comparing single answers against a reference or comparing multiple answers simultaneously, we decided to use it to compare generated CNs pairwise, as described in Section 4.1. This eliminates the problem of needing a reliable reference and instead focuses on determining which of the available options is the best, simplifying the task. Comparing the CNs against each other also avoids the ambiguity of the open-ended scenario we would face if we decided to evaluate them individually. We must note that JudgeLM also considers the instance of HS when comparing CNs, and thus evaluates CN effectiveness in relation to a specific instance of HS, instead as standalone sentences [^2] Appendix C.5 shows the template used to prompt JudgeLM..
JudgeLM operates in two modes: normal mode and fast evaluation mode. In fast evaluation mode the model outputs two scores, one for each CN, providing an overall assessment of their value. In normal mode the model supplements these scores with arguments explaining the rationale behind them. We used the normal mode both during the development stage and while conducting the result analysis, to ensure a comprehensive evaluation. However, we activated fast mode when creating the ranks, which are thus calculated based exclusively on the output scores. This decision was made because generating arguments significantly increases inference time and argumentation was deemed unnecessary for the ranking task.
@@ 4 Evaluation Framework
In this section we present a cost-effective pairwise rank-based evaluation paradigm designed to assess the performance of CN generation systems in alignment with human preference. The method is detailed in Section 4.1. In addition, in Section 4.2, we describe the additional manual evaluation performed to provide a detailed assessment according to various relevant aspects contributing to the effectiveness of a CN, as presented in (Bengoetxea et al., 2024).
@@ 4.1 Pairwise Rank-Based Evaluation
We propose an "A vs B" comparative setup to rank models with respect to their CN generation skills. Suppose we have $n$ models to rank and we want to evaluate their performance in a set consisting of $h$ HS instances. We first generate $h$ CNs, one for each HS instance, with each of the $n$ models. After that, the generated $h\cdot n$ CNs are pitted against each other in "A vs B" tournaments, with a total of (${n\choose 2}\cdot h=\frac{n!}{2!(n-2)!}\cdot h$ tournaments. Each tournament therefore comprises one HS and two CNs generated from different models, and is evaluated (either by automatic or manual methods) to assess which of the CN better answers the given HS. The model that produced the best CN receives $1$ point, while the losing model receives 00 points. In the case of ties, both models receive $0.5$ points. Models receive a total score by aggregating the scores of the tournaments they participate, and they are finally ranked based on the obtained score. In the following, we describe details of the manual and automatic evaluation methods, respectively.
@@
For automatic evaluation we prompt JudgeLM to output 2 scores: one for each proposed CN. The winner is determined by comparing these scores: the CN with the highest score is the winner. If both scores are the same, it results in a tie.